Grade 12General Physics

The current passing through a choke coil of 5 henry is decreasing at the rate of 2 amperejsec. The e.m.f. developing across the coil is
1. 10 V
2. 2.5 V
3. -10 V
4. -2.5 V

please explain whatever ans you give

kanika sharma
11 Years agoGrade 12

1 Answer

saurabh charde
11 Years ago
ANSWER IS 10 V.
e.m.f =−L×dI/dt
WHERE  L IS INDUCTANCE OF COIL AND dI/dt IS RATE OF CHANGE OF CURRENT
GIVEN: L = 5 henry  , dI/dt = −2 (minus sign because current is decreasing)
SO, e.m.f. = −5×(−2) = 10 V